# Simultaneous Ground Reaction Force and State Estimation via Constrained Moving Horizon Estimation

### v1

@@ IV-C Measurement Model
Leg Odometry: Once static contact is established between the foot and the ground, floating base velocity $y_{v}$ is measured based on the known leg kinematics, relying on the encoder position and velocity measurements $\tilde{\alpha}$ and $\tilde{\dot{\alpha}}$:
@@
$\tilde{y}_{v}$ $=-\hat{R}_{\text{WB}}J_{b}(\tilde{\alpha})\tilde{\dot{\alpha}}-\hat{R}_{\text{WB}}\tilde{\omega}^{\times}fk(\tilde{\alpha}),$ (20)
@@
Generalized Momentum: With known robot dynamics and encoder measurements, the generalized momentum measurement model can be formulated by similarly decoupling $M(q)$ as follows:
$M=\begin{bmatrix}M_{1}&M_{2}\end{bmatrix},$ (21)
where $M_{1}\in\mathbb{R}^{(n+6)\times 3}$ and $M_{2}\in\mathbb{R}^{(n+6)\times(n+3)}$. The generalized momentum measurement is then expressed as:
@@
$\tilde{y}_{m}$ $=M_{2}\begin{bmatrix}\tilde{\omega}\\ \tilde{\dot{\alpha}}\end{bmatrix},$ (23)
where $\delta^{y}_{m}\sim\mathcal{N}(0,R_{m})$ contains the measurement noise of $\omega$, $\alpha$ and $\dot{\alpha}$.
Visual Odometry: Visual Odometry (VO) estimates the robot’s pose in the camera frame $C$ by tracking features in the images captured by onboard cameras. Typically, the VO output is interpreted as the incremental homogeneous transformation between consecutive camera frames $C_{i}$ and $C_{j}$:
$\mathbf{T}_{\text{C}ij}=\mathbf{T}_{\text{WC}_{i}}^{-1}\mathbf{T}_{\text{WC}_{j}}+\delta_{\text{vo}},$ (24)
where $\mathbf{T}_{\text{WC}_{i}}$ and $\mathbf{T}_{\text{WC}_{j}}$ are the homogeneous transformations from the world frame $W$ to the camera frame $C$ at time $i$ and $j$. The VO measurement is corrupted by noise $\delta_{\text{vo}}\sim\mathcal{N}(0,Q_{\text{vo}})$. With the transformation $\mathbf{T}_{\text{BC}}$ between IMU/body frame $B$ and camera frame $C$ calibrated offline, the incremental transformation $\mathbf{T}_{\text{C}_{ij}}$ can be represented in the body frame as:
$\tilde{\mathbf{T}}_{\text{B}_{ij}}=\mathbf{T}_{\text{BC}}\tilde{\mathbf{T}}_{\text{C}_{ij}}\mathbf{T}_{\text{BC}}^{-1}.$ (25)
The output frequency of the VO matches the camera frame rate, ranging from 10 to 60 Hz. In our implementation, each camera/VO frame is aligned to the closest IMU frame on the time axis. After the alignment, only the translation component $p_{\text{B}_{ij}}$ of the incremental transformation $\mathbf{T}_{\text{B}_{ij}}$ is applied as a measurement into the estimation. We employ a Cubic Bézier polynomial to fit a series of relative translation measurements $\tilde{p}_{\text{B}_{ij}}$. A smooth path of position is generated using consecutive translations $\tilde{p}_{\text{B}_{ij}}$ as control points, providing interpolated VO measurements $\tilde{p}_{c_{k}}$ between consecutive IMU frames $k$ and $k+1$, $\forall k\in\{k\in\mathbb{N}^{+}|\ i\leq k\leq j-1\}$:
$y_{c_{k}}$ $=\hat{R}_{\text{WB}}^{\intercal}(p_{k+1}-p_{k})+\delta^{y}_{c_{k}},$ (26)
$\tilde{y}_{c_{k}}$ $=\tilde{p}_{c_{k}},$ (27)
@@ IV-D Physical Constraints
In [13], the Linear Complementary Problem (LCP) was formulated with the maximum dissipation principle and a polyhedral friction cone, leading to the development of a constrained Kalman filter for handling rigid slippery contacts. In this work, we utilize a simplified LCP to incorporate physical information into the constrained robot dynamics involved estimation. For static rigid contact, the contact complementarity constraint is formulated as:
$v_{\text{foot}}\tilde{b}$ $=0,$ (28)
$f(1-\tilde{b})$ $=0,$ (29)
$f_{\text{normal}}$ $\geq 0,$ (30)
where $v_{\text{foot}}$ and $f_{\text{normal}}$ denotes the foot velocity and normal component of the 3-DOF GRF on the corresponding robot foot. This simplification leverages the advantage of using a pressure sensor or contact switch as a contact indicator $\tilde{b}$, addressing the computational challenges introduced by the original mixed-integer formulation [13]. When walking on slippery terrain, (28) is reduced to $v_{\text{foot}}^{n}\tilde{b}=0$.
@@ IV-E Linear Moving Horizon Estimation
With linear constraints constructed above, the IV-E is formulated as a constrained Quadratic Program:
$\min_{\mathbf{X}_{[T-N,T]}}\ \Gamma(x_{T-N})$
$+\sum_{k=T-N}^{T-1}||\delta^{x}_{k}||_{Q_{k}}+\sum_{k=T-N}^{T}||\delta^{y}_{k}||_{R_{k}},$
@@ V Evaluation
Figure 2: (a) Sensor configuration of STRIDE. (b) Sensor configuration and experimental Setup of Unitree Go1.
Now we evaluate the proposed simultaneous force and state estimation method in both simulation and hardware, using the educational planar bipedal robot STRIDE [15] and commercial quadrupedal robot Unitree Go1, as illustrated in Fig. 2. The MHE is implemented using C++ in ROS2 and Ubuntu 20.04. The QP problem is solved using the off-the-shelf solver OSQP [28]. The VO is implemented through the open-source package ORB-SLAM3 [29] to perform VO at 60Hz. The MHEs are solved stably at 200Hz, with a window size of 8 that contains 1-2 VO frames in practice. All state estimations are performed in real-time on a PC with Intel Core i9-13900 CPU. Additionally, the STRIDE robot is controlled using S2S controller [24], and the Go1 robot is controlled using Model Predictive Control (MPC) [30] and Whole body control (WBC) [31] framework.
Table I: Sensor Noise of STRIDE and Unitree Go1 in simulation.
@@
[…]
We focus on comparing the GRF estimation errors of a modified Disturbance Kalman Filter (DKF) [11], Momentum Based Observer (MBO) [9], the proposed MHE method without physical constraints described in Section IV.D, and the full proposed MHE method. The modified DKF is formulated to be equivalent to an unconstrained MHE with a window size of one. For the MBO, we use the ground truth floating-base velocity to construct the generalized momentum, creating an idealized scenario for comparison.
V-A STRIDE Simulation
We evaluate our estimation method using simulation data from the educational planar bipedal robot STRIDE [15]. Fig. 3 and Table. II show our methods closely matched the ground truth, demonstrating that the proposed approach outperforms the MBO and DKF methods. Additionally, the inclusion of complementarity constraints, which enforce known physical information, further enhances the accuracy of the estimation.
Table III: RMSEs of the estimations on Go1 in simulation.
@@ V-A STRIDE Simulation
V-B Unitree Go1 Simulation
We evaluate our estimation method using simulation data from the Go1 quadruped robot. In simulation, the Go1 robot is controlled with a basic velocity command to execute omnidirectional locomotion with various gaits. Fig. 4 and Table. III shows the estimation results of floating-base velocity and GRF on the front left foot. The proposed method consistently demonstrates superior performance compared to canonical approaches.
Figure 4: Estimation results on of the GRF on the Unitree Go1 in simulation, $f_{z}$ and $f_{x}$, $f_{y}$, and $f_{z}$ denote the tangential frictional force and the vertical normal force, respectively, expressed in the world frame.
@@ V-C Unitree Go1 Hardware
Table IV: RMSEs of the estimations on Go1 hardware.
@@
We finally demonstrate our estimation method on the Go1 hardware. We obtain the stereo images from the Intel Realsense Depth Camera D455 and IMU data from the default Unitree IMU. During the experiments, the Go1 robot is controlled to trot on the force plate, as shown in Fig. 2. The comparison results on the estimation and ground truth of the GRF on the front left foot and base-linear velocity are shown in Fig. 5. As indicated in Table. IV, for the noisy hardware implementation, the constraint information effectively enforced the feasibility of the estimation results, leading to superior performance, even though the momentum-based methods were configured to use the ground truth floating-base velocity from the opti-track system. The inclusion of exteroceptive sensors improved the estimation during the stance phase, while the constraints further ensured the feasibility of the solution.
@@ VI Conclusions and Future Work
In the future, we aim to extend the proposed framework, exploring terrain dynamics by leveraging estimated foot position, velocity, and ground reaction forces to perform planning and control tasks on deformable and unstructured terrains. We will also focus on the computational algorithm design to improve the estimation frequency and window size and realize optimal calibrations on the process and measurement noises. Additionally, we are interested in integrating the Linear Complementarity Problem (LCP) within the windowed estimation framework to enable robust estimation in environments with rigid and slippery contact conditions.

### v2

@@ IV-C Measurement Model
Leg Odometry: Once static contact is established between the foot and the ground, floating-base velocity $y_{v}$ is measured based on the known leg kinematics which relies on the encoder position and velocity measurements $\tilde{\alpha}$ and $\tilde{\dot{\alpha}}$:
@@
$\tilde{y}_{v}$ $=-\hat{R}_{\mathcal{WB}}J_{b}(\tilde{\alpha})\tilde{\dot{\alpha}}-\hat{R}_{\mathcal{WB}}\tilde{\omega}^{\times}fk(\tilde{\alpha}),$ (16)
@@
Generalized Momentum: Given the robot’s dynamics and encoder measurements, the generalized momentum measurement model is formulated as:
@@
$\tilde{y}_{m}$ $=M_{2}\begin{bmatrix}\tilde{\omega}^{\intercal}&\tilde{\dot{\alpha}}^{\intercal}\end{bmatrix}^{\intercal},$ (18)
where $M_{1}\in\mathbb{R}^{(n+6)\times 3}$ and $M_{2}\in\mathbb{R}^{(n+6)\times(n+3)}$ are subblock of inertia matrix $M=\begin{bmatrix}M_{1}&M_{2}\end{bmatrix}$, and $\delta^{y}_{m}\sim\mathcal{N}(0,R_{m})$ contains the measurement noise of $\omega$, $\alpha$ and $\dot{\alpha}$.
Visual Odometry: Visual Odometry (VO) estimates the robot’s pose in the camera frame $\mathcal{C}$ by tracking features in images captured by onboard cameras. Typically, the VO output is interpreted as the incremental homogeneous transformation between consecutive camera frames $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$:
$\mathbf{T}_{\mathcal{C}ij}=\mathbf{T}_{\mathcal{WC}_{i}}^{-1}\mathbf{T}_{\mathcal{WC}_{j}}+\delta_{\text{vo}},$ (19)
where $\mathbf{T}_{\mathcal{WC}_{i}}$ and $\mathbf{T}_{\mathcal{WC}_{j}}$ are the homogeneous transformations from the world frame $\mathcal{W}$ to the camera frame $\mathcal{C}$ at time $i$ and $j$. The VO measurement is corrupted by noise $\delta_{\text{vo}}\sim\mathcal{N}(0,Q_{\text{vo}})$. With the transformation $\mathbf{T}_{\mathcal{BC}}$ between IMU/body frame $\mathcal{B}$ and camera frame $\mathcal{C}$ calibrated offline, the incremental transformation $\mathbf{T}_{\mathcal{C}_{ij}}$ can be represented in the body frame as:
$\tilde{\mathbf{T}}_{\mathcal{B}_{ij}}=\mathbf{T}_{\mathcal{BC}}\tilde{\mathbf{T}}_{\mathcal{C}_{ij}}\mathbf{T}_{\mathcal{BC}}^{-1}.$ (20)
The output frequency of the VO matches the camera frame rate, ranging from 10 to 60 Hz. In our implementation, each camera/VO frame is aligned to the closest IMU frame on the time axis. After the alignment, only the translation component $p_{\mathcal{B}_{ij}}$ of the incremental transformation $\mathbf{T}_{\mathcal{B}_{ij}}$ is applied as a measurement into the estimation. We employ a Cubic Bézier polynomial to fit a series of relative translation measurements $\tilde{p}_{\mathcal{B}_{ij}}$. A smooth path of position is generated using consecutive translations $\tilde{p}_{\mathcal{B}_{ij}}$ as control points, providing interpolated VO measurements $\tilde{p}_{c_{k}}$ between consecutive IMU frames $k$ and $k+1$, $\forall k\in\{k\in\mathbb{N}^{+}|\ i\leq k\leq j-1\}$:
$y_{c_{k}}=\hat{R}_{\mathcal{WB}}^{\intercal}(p_{k+1}-p_{k})+\delta^{y}_{c_{k}},\quad\tilde{y}_{c_{k}}=\tilde{p}_{c_{k}},\$ (21)
@@ IV-D Physical Constraints
In [13], the Linear Complementary Problem (LCP) is formulated with the maximum dissipation principle and a polyhedral friction cone, leading to the development of a constrained Kalman Filter for handling rigid slippery contacts. In this work, we employ a simplified LCP to integrate physical information as constraints in the estimation. For static rigid contact, the complementarity constraint is formulated as:
$v_{\text{foot}}\tilde{b}$ $=0,\quad f(1-\tilde{b})=0,\quad f_{\text{normal}}\geq 0,$ (22)
where $v_{\text{foot}}$ and $f_{\text{normal}}$ denote the foot velocity and the normal component of the linear part of GRF acting on the corresponding robot foot, respectively. This simplification leverages the advantage of using a pressure sensor or contact switch as contact indicator $\tilde{b}$, addressing the computational challenges introduced by the original mixed-integer formulation [13]. When walking on slippery terrain, the velocity constraint of the contact foot reduces to $v_{\text{foot}}^{n}\tilde{b}=0$.
@@ IV-E Linear Moving Horizon Estimation
With linear constraints constructed above, the (III-A) is formulated as a constrained Quadratic Program:
$\min_{\mathbf{X}_{[T-N,T]}}\ \Gamma(x_{T-N})\textstyle$
$+\textstyle{\sum}_{k=T-N}^{T-1}||\delta^{x}_{k}||_{Q_{k}}+\textstyle{\sum}_{k=T-N}^{T}||\delta^{y}_{k}||_{R_{k}},$ (23)
@@
Figure 2: Sensor configurations of (a) Bucky and (b) STRIDE. (c) Sensor configuration and experimental setup of the Unitree Go1.
@@ V Evaluation
Now we evaluate the proposed simultaneous force and state estimation method in both simulation and hardware, using the custom-designed humanoid robot Bucky, the educational planar bipedal robot STRIDE [15] and commercial quadrupedal robot Unitree Go1, as illustrated in Fig. 2. The MHE is implemented using C++ in ROS2 and Ubuntu 20.04. The QP problem is solved using the off-the-shelf solver OSQP [28]. The VO is implemented through the open-source package ORB-SLAM3 [29] to perform VO at 60Hz. The MHEs are solved stably at 200Hz, with a window size of 8 that contains 1-2 VO frames in practice. All state estimations are performed in real-time on a PC with Intel Core i9-13900 CPU. Additionally, the STRIDE and Bucky robots are controlled using the S2S controller [24], [30], and the Go1 robot is controlled using Model Predictive Control (MPC) and Whole body control (WBC) framework [31], [32], [33].
Table I: Sensor Noise of Bucky, STRIDE and Go1 in simulation.
@@
Table III: RMSEs of the estimations on Bucky in simulation.
Estimation Our Method Our Method DKF MBO
Method (w/o IV.D)
$\text{RMSE}_{v}$ [m/s] 0.0161 0.0165 0.0153 NA
$\text{RMSE}_{f}$ [N] 10.4337 24.6590 33.1068 10.1655
Figure 5: Estimation results of the GRF on Bucky in simulation, $f_{x}$, $f_{y}$, and $f_{z}$ denote the tangential frictional force and the vertical normal force, respectively, expressed in the world frame.
The simulation is configured with the MuJoCo [34], which runs the time-stepping simulation at 500 Hz. The sensor noise settings for the MuJoCo environment are detailed in Table. I, and the ground truth GRF is obtained from a force sensor attached to the foot. For hardware experiments, we use a motion-capture system that has 12 OptiTrack Prime-13 cameras [35] to obtain the ground truth position and orientation at 200Hz. The ground truth velocity is obtained by computing the central differences of the position. The GRF ground truth is obtained from a Bertec-4060 6D Force Plate [36] at 1000Hz, with a resolution of ±$\pm 0.4\text{N}$. Fig. 3 illustrates the system architecture for estimation.
We focus on comparing the GRF estimation errors of a modified Disturbance Kalman Filter (DKF) [11], Momentum-Based Observer (MBO) [9], the proposed MHE method without physical constraints described in Section IV.D, and the full proposed MHE method. The modified DKF is formulated to be equivalent to an unconstrained MHE with a window size of one. For the MBO, we use the ground truth floating-base velocity to construct the generalized momentum, creating an idealized scenario for comparison.
V-A Humanoid Bucky and Bipedal STRIDE Simulations
We evaluate our estimation method using simulation data from the humanoid robot Bucky and planar bipedal robot STRIDE [15]. Fig. 4, Fig. 5, Table. II, and Table. III show our methods closely match the ground truth, demonstrating that the proposed approach outperforms the MBO and DKF methods. Additionally, the inclusion of contact complementarity constraints, which enforces known physical information, further enhances the accuracy of the estimation.
V-B Unitree Go1 Simulation
We evaluate our method using simulation data from the Go1 quadruped robot. In simulation, the Go1 robot is controlled with a basic velocity command to execute omnidirectional locomotion with various gaits. Fig. 6 and Table. IV show the estimation results of floating-base velocity and GRF on the front left foot. The proposed method demonstrates superior performance compared to canonical approaches.
Table IV: RMSEs of the estimations on Go1 in simulation.
@@ V-B Unitree Go1 Simulation
Figure 6: Estimation results of the GRF on the Unitree Go1 in simulation.
@@ V-C Unitree Go1 Hardware
Table V: RMSEs of the estimations on Go1 hardware.
@@
We finally demonstrate our estimation method on the Go1 hardware. We obtain the stereo images from the Intel Realsense Depth Camera D455 and IMU data from the default Unitree IMU. During the experiments, the Go1 robot is controlled to trot on the force plate, as shown in Fig. 2. The comparison results between the estimation and ground truth of the GRF on the front left foot and base-linear velocity are shown in Fig. 7. As indicated in Table. V, for the noisy hardware implementation, the constraint information effectively enforces the feasibility of the estimation results, leading to superior performance, even though the momentum-based methods are configured to use the ground truth floating-base velocity from the opti-track system. The inclusion of exteroceptive sensors improves the estimation during the stance phase, while the contact constraints further ensure the feasibility of the solution.
@@ VI Conclusions and Future Work
In the future, we aim to explore terrain dynamics by leveraging estimated foot position, velocity, and GRF to perform planning and control tasks on deformable and unstructured terrains. We will also focus on designing computational algorithms to improve estimation frequency and window size while optimizing the calibration of process and measurement noises. Additionally, we are interested in integrating the Linear Complementarity Problem (LCP) within the windowed estimation framework to enable robust estimation in environments with rigid and slippery contact conditions.